# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

### v1

@@ 5 Numerical results
Figure 1: Scan samples under successive constraints in the ($\mu_{\gamma\gamma}$, $\mu_{\tau\tau}$) panel with color coded by $\chi^{2}_{\rm 95}$. The left panel applies perturbative unitarity, the positive-definiteness vacuum-stability conditions, and passes the HiggsBounds/HiggsSignals tests. The right panel additionally imposes constraints from $B$-physics, the oblique parameters $(S,T,U)$, and the Higgs $\kappa$ fits.
Based on the scanned samples, the single light custodial singlet $H$ in the GM model near 95.4 GeV can account for the LHC $\gamma\gamma$ and $\tau\tau$ excesses at the $2\sigma$ level. Quantitatively, the scan covers the reported $2\sigma$ region in $\mu_{\gamma\gamma}$, while $\mu_{\tau\tau}$ typically peaks near 0.55 after all cuts. This underscores the lingering tension discussed at the end of Sec.  3 and leads to a definitive result.
Figure 1 shows the scan in the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane, with the color indicating $\chi^{2}_{95}$ defined in Eq. (41). The left panel applies perturbative unitarity, the positive-definiteness vacuum-stability conditions, and the HiggsBounds/HiggsSignals constraints. The right panel further incorporates the $B$-physics observations, electroweak precision constraints $(S,T,U)$, and fits to the SM-like Higgs $\kappa$ parameters. In the left panel, the samples are distributed almost uniformly across $\mu_{\gamma\gamma}\in[0.08,0.40]$, while $\mu_{\tau\tau}$ shows a different pattern. The lowest $\chi^{2}_{95}$ values are concentrated along a ridge where moderate values of $\mu_{\gamma\gamma}$ (approximately 0.20 to 0.3) coincide with $\mu_{\tau\tau}$ values around 0.50 to 0.54. The maximum value of $\mu_{\tau\tau}$ can reach up to 0.61. After incorporating $B$ physics, $(S,T,U)$, and the $\kappa$-constraints (right panel), the distribution becomes more compact: the high-$\mu_{\tau\tau}$ tail above approximately 0.55 and the extreme edges of $\mu_{\gamma\gamma}$ are eliminated. This behavior reflects the interplay between the charged-scalar loop enhancement of $H\to\gamma\gamma$ and the rescaling of fermionic couplings, which constrains $\mu_{\tau\tau}$ as the total width increases, as we discussed in Sec. 3.
Figure 2: Similar to Fig. 1: Top panel shows the ($\kappa_{ff}$, $\kappa_{VV}$) plane, with color indicating the $p$-value as calculated by HiggsSignals. The middle panel displays the ($\alpha$, $\mathcal{B}(B\to X_{s}\gamma)$) plane, with color representing $\mathcal{B}(B_{s}^{0}\to\mu^{+}\mu^{-})$. The bottom panel presents the distributions of the oblique parameters $S$, $T$, and $U$.
In Fig. 2, we present the distributions of the effective Higgs couplings $\kappa_{ff}^{h}$ and $\kappa_{VV}^{h}$ for the SM-like Higgs state $h$, the singlet mixing angle $\alpha$, and the oblique parameters $S$, $T$, and $U$ to illustrate their effects. Before applying these constraints, it is evident that the samples distributed into three distinct clusters. All three clusters require significant mixing between the doublet state $\phi$ and the singlet from the $\Delta$ field. Consequently, the mixing angle $\alpha$ must be approximately $\frac{\pi}{3}$ or $\frac{2\pi}{3}$, both of which correspond to similar values of $\left|\kappa_{ff}^{H}\right|^{2}$. In this case, the lighter $CP$-even Higgs state $H$ can acquire a sufficiently large Yukawa coupling, which in turn significantly affects its $ggF$ production rate and total decay width.
The first type of sample is characterised by $\alpha\approx 1.05$ ($\approx\frac{\pi}{3}$). For this type samples, although $\Gamma(H\to\tau\tau)$ is relatively large, the total width $\Gamma^{\rm tot}_{H}$ does not increase significantly for two reasons. First, with $m_{H}$ at only about 95 GeV, many decay channels—such as $t\bar{t}$, $WW$, and ZZZZ—are either kinematically forbidden or highly suppressed. Second, when $\alpha\approx 1$, the coupling of $H$ to gauge bosons is not particularly strong, so the total width does not increase substantially due to $VV$ modes. As a result, the denominator $\Gamma^{\rm tot}_{H}$ does not become large enough to "dilute" the enhancement in the numerator. Consequently, $\mu_{\tau\tau}$ can remain around $\sim 0.6$, which is higher than in the case with $\alpha\approx 2.1$. In contrast, in the branch with $\alpha\approx 2.1$, the gauge boson couplings of $H$ are closer to the SM structure. As a result, the contributions from $H\to VV$ and other lepton/quark channels to the total width are larger, leading to an increase in $\Gamma^{\rm tot}_{H}$ and consequently a suppression of $\mu_{\tau\tau}$. This highlights the tension between the SM-like Higgs state fit and the interpretation of the 95 GeV signal within the GM model. The first type of samples also predict a lower value of $\kappa^{h}_{VV}$ and are therefore completely excluded by Higgs effective coupling measurements, as shown in the top panel of Fig. 2.
Figure 3: Similar to Fig. 1: Top panel shows the ($m_{H},v_{\Delta}$) plane, with color indicating the $p_{\text{value}}$; bottom panel shows the ($m_{h},m_{H_{5}}$) plane, with color representing $m_{H_{3}}$.
To further illustrate the remaining two classes with $\alpha\sim 2.1$, we present plots of the samples in the ($m_{H}$, $v_{\Delta}$) plane, color-coded by $p$-value, and in the ($m_{h}$, $m_{H_{5}}$) plane, with colors representing $m_{H_{3}}$, as shown in Fig. 3. The second type of samples is characterized by $v_{\Delta}$ greater than 10 GeV, while the remaining samples have $v_{\Delta}$ less than 10 GeV. In the GM model, for the 125 GeV Higgs to exhibit SM-like behavior, its effective couplings to gauge bosons ($W$, $Z$) and fermions ($t$, $b$, $\tau$) must closely match their SM values. This requirement is directly linked to the doublet-triplet mixing angle $\theta$ and the triplet vacuum expectation value (VEV) $v_{\Delta}$. When $v_{\Delta}$ is small ($\lesssim 3$ GeV), the triplet component in the SM-like Higgs state $h$ is negligible, resulting in a significantly reduced coupling of the 125 GeV Higgs to $W$ and $Z$ bosons ($\kappa_{VV}^{h}<1$). This deviation undermines the SM-like behavior of the Higgs and creates tension with global $\kappa$ fits from the LHC. Conversely, when $v_{\Delta}$ is large ($\gtrsim 10$ GeV), the gauge boson coupling of the 125 GeV Higgs (see Eq. (22)) can be tuned to be very close to 1. In this regime, the decay and production rates of $h$ align with SM predictions, satisfying the constraints from HiggsSignals. In other words, a larger $v_{\Delta}$ (>10>10 GeV) can effectively "absorb" the effects of new physics, making the 125 GeV Higgs $h$ more SM-like. This trend is evident in numerical scans: when $v_{\Delta}$ falls within the [10, 15] GeV range, the surviving parameter points yield higher $p$-values (as reported by HiggsSignals), indicating better agreement with experimental data.
From the bottom panels of Fig. 2, The distributions of the oblique Peskin-Takeuchi parameters $T$, $S$, and $U$ differ significantly among the three sample classes. This is fairly straightforward. The $T$ parameter is often regarded as a sensitive probe of custodial symmetry breaking. In the GM model, custodial symmetry is preserved at tree level, so the primary contributions to $T$ come from mass splittings within the triplet and quintuplet states. For all samples, the custodial symmetry is nearly exact, resulting in $T\approx 0$. For larger $v_{\Delta}$ or varying $\alpha$, mixing between the light scalar $H$ and the 125 GeV Higgs can shift the masses of certain scalars up or down, leading to mass differences within custodial multiplets and thus deviations in $T$. The $S$ parameter represents corrections to the $Z$ boson self-energy and its mixing, arising from new physics states. In GM model, $S$ parameter is sensitive to light charged scalars ($H_{3}^{\pm}$, $H_{5}^{\pm}$, $H_{5}^{\pm\pm}$), which shown in Fig 3. When the light scalar $H$ has a larger triplet component, for example in the region $\alpha\approx 1$, loop effects are enhanced, leading to more significant deviations in $S$. Conversely, if the 125 GeV Higgs remains SM-like, for example $\alpha\simeq 2.1$ and large $v_{\Delta}$, the new states are heavier or have smaller mixing, resulting in smaller corrections to $S$.
Let us turn to the $B$-rare decays. The $B$-rare decay processes are highly suppressed in the SM, dominated mainly by top quark and $W$ boson loops. While in the GM model, new charged scalars such as $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ are introduced, which can also participate in these loop diagrams. The masses of these charged scalars and their couplings to fermions (which depend on $\alpha$ and $v_{\Delta}$) vary across different parameter regions, leading to different corrections to the $B$-rare decay amplitudes. In the so-called "alignment limit" of the GM model, the constraints from rare B-meson decays are not very stringent. As shown in these two panels, all sample points fall within the allowed range. Therefore, as discussed earlier, although the branching ratios for rare $B$ decays are of similar magnitude, they can still be classified into three distinct groups.
Finally, the large signal strength observed in the di-tau excess appears somewhat inconsistent with recent experimental constraints from searches for Higgs boson production in association with either $t\bar{t}$ or a $Z$ boson, followed by decay into tau pairs, as reported by CMS CMS:2024ulc . This tension is further heightened by LEP results from searches for the process $e^{+}e^{-}\to Z(\phi\to\tau^{+}\tau^{-})$ LEP:2003ing . Although the observed value is still compatible with current experimental limits, it will be thoroughly tested at the HL-LHC ZurbanoFernandez:2020cco ; LHCb:2018roe (with an expected precision of $3.4\%$ for $\kappa_{t}$) and at CEPC CEPCPhysicsStudyGroup:2022uwl ; CEPCStudyGroup:2023quu ; Ai:2024nmn ; Antusch:2025lpm (with an expected precision of less than $1\%$ for $\kappa_{t}$). Recent global fits of GM models further support this conclusion, indicating that the realistic signal strength in the di-tau channel should be significantly lower. It is also worth noting that the ATLAS collaboration has not yet published results from searches in the di-tau final state for masses around 95 GeV. Forthcoming results from ATLAS are therefore expected to clarify the situation, potentially reducing the central value of the $\tau\tau$ excess and providing a clearer understanding of the underlying physics.

### v2

@@ 5 Numerical results
In this section, we present detailed results from our sampling. For clarity, we have filtered out all samples that fail the HiggsBounds or HiggsSignals consistency requirements, the positive-definiteness requirement, or the perturbative unitarity constraints; thus, all samples shown in the figures satisfy $\Theta_{\tt HB}=1$, $\Theta_{\tt HS}=1$, $\Theta_{\rm PD}=1$, and $\Theta_{\rm unit}=1$.
5.1 Fit results and benchmarks
Figure 1: The profile likelihood (PL) map projected onto the plane of signal strengths for di-photon and di-tau channels, $(\mu_{\gamma\gamma},\mu_{\tau\tau})$.
Figure 1 displays the PL distribution projected onto the $(\mu_{\gamma\gamma},\,\mu_{\tau\tau})$ plane for the $95~\mathrm{GeV}$ scalar in the GM model, after applying all theoretical and experimental constraints described in Sec. 4. The color scale indicates the PL normalized to its global maximum. A well-defined likelihood concentration is observed in our fit, indicating that the GM model is capable of simultaneously enhancing the di-photon and di-tau signal strengths relative to the SM expectations. The highest-likelihood region is centered around $\mu_{\gamma\gamma}\sim 0.25$ and $\mu_{\tau\tau}\sim 0.5$, demonstrating that the reported di-photon excess can be accommodated within the preferred parameter region, while the di-tau channel is enhanced to a moderate but non-negligible level.
Refer to caption Refer to caption Refer to caption
Figure 2: Left: The PL map on the $(\kappa_{F},\kappa_{V})$ plane. Middle: The PL $\mathcal{L}_{\kappa}$ on the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane. Right: The PL $\mathcal{L}_{\rm tot}$ without $\mathcal{L}_{\kappa}$ contribution on the $(\mu_{\gamma\gamma},\mu_{\tau\tau})$ plane.
In summary, the main difficulty in accommodating the 95 GeV $\gamma\gamma$ and $\tau\tau$ excesses within the GM model originates from the tension with the precision $\kappa$ fits of the 125 GeV Higgs couplings, which restrict the viable region to a thin and highly correlated slice of parameter space. As shown in Eq. (24), the signal strength $\mu_{\tau\tau}$ is directly related to $\kappa_{F}$. To illustrate this tension between the excesses and the Higgs-coupling data, Fig. 2 shows the PL distribution on the $(\kappa_{V},\kappa_{F})$ plane, the profile behavior of $\mathcal{L}_{\kappa}\equiv\exp\left(-\frac{1}{2}\chi^{2}_{\kappa}\right)$, and $\mathcal{L}_{\rm tot}/\mathcal{L}_{\kappa}$. The left panel of Fig. 2 shows a striking feature of this distribution: the emergence of a narrow, elongated band with a clear positive correlation between $\kappa_{F}$ and $\kappa_{V}$. This behavior can be directly understood from the analytic expressions in Eq. (22), which show that both coupling modifiers are controlled solely by the two mixing angles $\alpha$ and $\sin{\theta_{H}}$. $\kappa_{F}$ and $\kappa_{V}$ are strongly correlated through the underlying doublet-triplet mixing structure, and their allowed values populate only a restricted trajectory in the $(\kappa_{F},\kappa_{V})$ plane. As a consequence, the GM model can only access a limited "half-ellipse" structure observed in the PL map. In the parameter region preferred by the fit, an increase in $|\cos\alpha|$ simultaneously enhances both $\kappa_{F}$ and $\kappa_{V}$, leading to the positively sloped band observed in the figure. Moreover, a large $\kappa_{F}$ implies a sizable doublet component in the lighter 95 GeV state $H$, enhancing its gluon-fusion production cross section and leading to tension with direct searches encoded in HiggsBounds. These combined effects carve out the characteristic blank wedge in the distribution. As a consequence, in the middle panel of Fig. 2, the best $\mathcal{L}_{\kappa}$ is located in the region of $\mu_{\tau\tau}$ around 0.4. In the right panel of Fig. 2, the remaining contribution to $\mathcal{L}_{\rm tot}$ is larger in the region with higher $\mu_{\tau\tau}$.
Table 1: Benchmark points in the GM model. The $\chi^{2}_{\rm tot}$ is defined in Eq. (30). A checkmark (✓) and a cross (✗) indicate whether a given benchmark point passes or fails the corresponding constraint. All dimensional quantities are given in units of GeV.
BP1 BP2 BP3 BP1 BP2 BP3 BP1 BP2 BP3
$\lambda_{2}$ 0.8211-0.8211 0.6069-0.6069 $0.0698$ $m_{H}$ $96.8491$ $96.68$ $95.22$ $\mu_{\gamma\gamma}$ $0.2899$ $0.2368$ $0.2445$
$\lambda_{3}$ $0.8150$ $0.8336$ 0.8805-0.8805 $m_{h}$ $124.0331$ $125.2977$ $125.0191$ $\mu_{\tau\tau}$ $0.7423$ $0.5245$ $0.5445$
$\lambda_{4}$ $0.5232$ 0.0417-0.0417 $0.4704$ $m_{H_{3}}$ $194.5481$ $214.3225$ $222.6281$ $\chi^{2}_{\rm tot}$ $19.795$ $21.220$ $7.605$
$\lambda_{5}$ $0.2216$ $0.8814$ 0.0418-0.0418 $m_{H_{5}}$ $290.7393$ $340.3623$ $311.1941$ $\mathcal{B}(H_{5}^{\pm\pm}\!\to\!W^{\pm}W^{\pm})$ $0.1757$ $0.0032$ $0.1059$
$\alpha$ 0.8075-0.8075 $1.0548$ 1.0399-1.0399 $\kappa_{F}^{h}$ $0.8538$ $0.9328$ $0.9330$ $\mathcal{B}(H_{5}^{\pm\pm}\!\to\!H_{3}^{\pm}W^{\pm})$ $0.8243$ $0.9968$ $0.8941$
$\sin\theta_{H}$ $0.1309$ $0.0287$ $0.1246$ $\kappa_{V}^{h}$ $0.9296$ $0.9200$ $0.9836$ Unitarity
$M_{H}^{\rm in}$ $96.84$ $96.6780$ $95.2205$ $S$ 0.0832-0.0832 0.0930-0.0930 0.0850-0.0850 Positive definiteness
$M_{h}^{\rm in}$ $124.0331$ $125.2977$ $125.0191$ $T$ 0.0136-0.0136 0.0012-0.0012 0.0181-0.0181 HiggsSignals
$U$ $0.0007$ $0.0003$ $0.0004$ HiggsBounds
To highlight the characteristic features of this surviving region, we present three benchmark points in Table. 1. BP1 features a relatively large triplet admixture, with $\sin\theta_{H}=0.131$ and $\alpha=-0.808$. This results in a significant suppression of the SM-like Higgs coupling $\kappa_{F}$. Consequently, the 95 GeV scalar attains a sizable $\mu_{\tau\tau}=0.742$ and the highest diphoton rate among the three points, $\mu_{\gamma\gamma}=0.290$, leading to the smallest $\chi^{2}_{\rm 95}$. However, it does not satisfy the HiggsSignals constraints. BP2 represents the opposite limit with a very small triplet VEV, $\sin\theta_{H}=0.029$. The 125 GeV Higgs is therefore more SM-like. The reduced mixing suppresses the $\tau\tau$ signal strength to $\mu_{\tau\tau}=0.525$, while the diphoton rate, $\mu_{\gamma\gamma}=0.237$, is mainly maintained by charged-scalar loop effects. This benchmark satisfies both HiggsSignals and HiggsBounds. Compared to BP1, BP3 provides a better solution with $\sin\theta_{H}=0.125$ and $\alpha=-1.040$, for which $\kappa_{V}^{h}=0.984$ is remarkably close to experimental data. This point achieves a balanced configuration, $\mu_{\tau\tau}=0.545$ and $\mu_{\gamma\gamma}=0.245$, while remaining consistent with all theoretical and experimental constraints. It exemplifies the thin, correlated region identified in Fig. 2, where the 95 GeV excesses can be accommodated without significantly disturbing the 125 GeV Higgs couplings.
Figure 3: Similar to Fig. 1, but projected onto the GM model parameter planes.
5.2 Parameter space and phenomenology
Figure 3 displays the PL projected onto the eight GM model parameter planes. A clear pattern emerges: while the quartic couplings $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ span relatively broad ranges, the physically relevant mixing parameters are tightly constrained. In the two $\lambda_{i}$ panels, no pronounced minima appear, indicating that the quartic interactions are not individually strongly constrained. Instead, these couplings primarily support the scalar mass spectrum and enable viable loop contributions, without requiring fine-tuned correlations among themselves. In contrast, the $(M_{H}^{\rm in},\alpha)$ projection reveals a highly concentrated high-likelihood region centered at $(95~{\rm GeV},-\frac{\pi}{3})$, directly corresponding to the mass of the observed excess. The narrow range of the mixing angle $\alpha$ is consistent with the data, and large portions of the $\alpha$ parameter space are excluded, indicating that doublet–triplet mixing is tightly constrained. For values of $\alpha$ with identical magnitude as shown in the benchmark points BP2 and BP3, $\kappa_{F}$ remains unchanged, as expected. However, the coupling $\kappa_{V}$ associated with positive $\alpha$ is significantly smaller than that for negative $\alpha$, resulting in a pronounced deviation from experimental constraints. Moreover, $\kappa_{V}^{H}$ corresponding to positive $\alpha$ is substantially larger than its counterpart for negative $\alpha$, rendering this scenario unable to survive the additional Higgs-search constraints in the HiggsBounds package. This feature underlies the $\kappa$ tension observed in Fig. 2: while $\alpha$ at $\frac{\pi}{4}$ would better explain the excess, the overall fit only allows $\alpha$ near $\frac{\pi}{3}$. A similar pattern appears in the $(M_{h}^{\rm In},\sin\theta)$ plane. Overall, Fig. 3 demonstrates that the triplet admixture must remain moderate.
Figure 4: Similar to Fig. 1, but shown in the oblique-parameter $(T,S)$ plane and in the branching-ratio plane for the rare $B$-meson decays $B^{0}_{s}\to\mu^{+}\mu^{-}$ and $B\to X_{s}\gamma$.
The results for the branching ratios of rare $B$-meson decays and the EW precision observables $T$ and $S$ are shown in Fig. 4. These rare $B$-decay processes are highly suppressed in the SM and are dominated mainly by top-quark and $W$-boson loops. In the GM model, new charged scalars such as $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ also participate in these loop diagrams. In our results, the two branching ratios are centered around the experimental central values and exhibit a strong negative correlation. This anti-correlation can be understood at the level of the effective weak Hamiltonian. Both observables are loop-induced flavor-changing neutral-current processes and are therefore sensitive to the charged-scalar sector of the GM model, in particular to the custodial triplet state $H_{3}^{\pm}$. The radiative decay $B\to X_{s}\gamma$ is governed by the electromagnetic dipole operator
$\mathcal{H}_{\rm eff}\supset-\frac{4G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}C_{7}^{\rm eff}\mathcal{O}_{7},\quad\mathcal{O}_{7}=\frac{e}{16\pi^{2}}m_{b}(\bar{s}\sigma^{\mu\nu}P_{R}b)F_{\mu\nu}.$ (43)
The branching ratio scales approximately as $\mathcal{B}(B\to X_{s}\gamma)\propto|C_{7}^{\rm eff}|^{2}$ Buras:1998raa ; Misiak:2015xwa . In contrast, the rare decay $B_{s}^{0}\to\mu^{+}\mu^{-}$ is described by
$\mathcal{H}_{\rm eff}\supset-\frac{4G_{F}}{\sqrt{2}}V_{tb}V_{ts}^{*}\left(C_{10}\mathcal{O}_{10}+C_{S}\mathcal{O}_{S}+C_{P}\mathcal{O}_{P}\right),\quad\mathcal{O}_{10}=(\bar{s}\gamma^{\mu}P_{L}b)(\bar{\mu}\gamma_{\mu}\gamma_{5}\mu),$ (44)
where $\mathcal{O}_{S,P}$ denote scalar and pseudoscalar operators. The corresponding branching ratio depends predominantly on $|C_{10}|^{2}$ in the absence of large scalar contributions Bobeth:2013uxa ; Hermann:2013kca . In the GM model, the charged scalars $H_{3}^{\pm}$, $H_{5}^{\pm}$, and $H_{5}^{\pm\pm}$ modify both $C_{7}$ and $C_{10}$ at the one-loop level. However, the loop functions entering $\Delta C_{7}$ and $\Delta C_{10}$ differ in structure and interfere differently with the SM contributions. For the parameter region relevant to the 95 GeV interpretation, the charged-scalar contribution tends to shift $C_{7}$ and $C_{10}$ in opposite effective directions relative to their SM values, leading to the nearly linear anti-correlation observed in Fig. 4. Turning to the EW oblique measurements, the $T$ parameter is often considered a sensitive probe of custodial symmetry breaking. In the GM model, custodial symmetry is preserved at tree level, so the main contributions to $T$ arise from mass splittings within the triplet and quintuplet states. For all samples, custodial symmetry remains nearly exact, resulting in $T\approx 0$. However, for larger $v_{\Delta}$ or varying $\alpha$, mixing between the light scalar $H$ and the 125 GeV Higgs can shift the masses of certain scalars, causing mass differences within custodial multiplets and leading to deviations in $T$. The $S$ parameter reflects corrections to the $Z$ boson self-energy and its mixing, originating from new physics states. In the GM model, the $S$ parameter is sensitive to light charged scalars. When the light scalar $H$ has a larger triplet component, such as in the region $\alpha\sim-\pi/3$, loop effects are enhanced, resulting in more significant deviations in $S$. Conversely, when the 125 GeV Higgs boson is near the alignment limit, such as for $\alpha\sim\pi/3$ and a small triplet vacuum expectation value $v_{\Delta}$, the scalar sector is dominated by the doublet component, and mixing with the triplet states is suppressed. In this regime, the additional custodial multiplet scalars either become heavier or couple more weakly to the electroweak gauge bosons. As a result, their loop-induced contributions to the gauge-boson self-energies are reduced, leading to smaller corrections to the oblique parameter $S$.
Figure 5: Similar to Fig. 1, but showing the $(m_{H_{5}},m_{H_{3}})$ panel and the $m_{H_{5}^{\pm\pm}}$ versus $\sin\theta\times\sqrt{\mathcal{B}\left(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm}\right)}$ panel. The red line indicates the ATLAS search ATLAS:2023sua limit on doubly charged $H_{5}^{\pm\pm}$ in $W^{\pm}W^{\pm}$ final states.
Finally, Fig. 5 displays the profiled likelihood projected onto the $(M_{H_{5}},\,M_{H_{3}})$ plane. A compressed spectrum is observed between the custodial fiveplet and triplet masses, with $\Delta m(m_{H_{5}},m_{H_{3}})<\mathcal{O}(100~{\rm GeV})$. In particular, the robust mass hierarchy $M_{H_{5}}>M_{H_{3}}$ implies that the doubly charged scalar $H_{5}^{\pm\pm}$ tends to decay predominantly into $H_{3}^{\pm}W^{\pm}$, rather than through the conventional $W^{\pm}W^{\pm}$ mode. As shown in Table 1, the branching ratio $\mathcal{B}(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm})$ is typically suppressed to below the ten-percent level, and for the benchmark point BP2 it can even reach the per-mille level. This feature has important implications for collider phenomenology.
The right panel of Fig. 5 shows the plane of $m_{H_{5}^{\pm\pm}}$ versus $\sin\theta\,\sqrt{\mathcal{B}(H_{5}^{\pm\pm}\to W^{\pm}W^{\pm})}$, together with the published ATLAS limit ATLAS:2023sua derived for the simplified topology with dominant $H_{5}^{\pm\pm}\to W^{\pm}W^{\pm}$ decays. We emphasize that this comparison is only indicative. In the cascade-dominated regime relevant to our preferred parameter space, the actual signal topology is instead controlled by $H_{5}^{\pm\pm}\to H_{3}^{\pm}W^{\pm}$, and the resulting event kinematics and analysis efficiencies need not coincide with those assumed in the published same-sign diboson search. Therefore, the experimental contour in Fig. 5 should not be interpreted as a strict exclusion bound on the full GM scenario in this region.
Nevertheless, the figure still illustrates an important trend: once the branching ratio into $W^{\pm}W^{\pm}$ is strongly suppressed, the direct sensitivity of the existing same-sign diboson searches is correspondingly weakened. A dedicated recast of the relevant ATLAS and CMS analyses, explicitly incorporating the cascade topology $H_{5}^{\pm\pm}\to H_{3}^{\pm}W^{\pm}$, would be required to assess the true collider reach for this part of the parameter space.
Additionally, for the Drell–Yan process, which produces $H_{5}$ pairs via $\gamma$ and $Z$ exchange ATLAS:2021jol , the production cross section depends only on the gauge coupling and the Higgs mass $m_{H_{5}}$, and is independent of $v_{\Delta}$ as discussed in the experimental searches. In our work, $m_{H_{5}}$ ranges from several hundred GeV, where the cross section is relatively small, making it difficult for current experimental searches to exclude the surviving parameter space. The process $H_{5}^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$ proceeds via a lepton number violation coupling between the triplet $\chi$ and lepton fields, conducted at LEP OPAL:2001luy . However, as pointed out in Chiang:2012cn , this decay channel becomes comparable to the weak gauge boson decay modes only when $v_{\Delta}\lesssim 10^{-4}~\text{GeV}$, making it negligible in the parameter region of interest for our study.
5.3 Additional checks
As discussed in Sec. 4, our main fit uses the $(\kappa_{V},\kappa_{F})$ likelihood rather than the full HiggsSignals output, in order to avoid double counting Higgs signal-strength information. It is nevertheless useful to verify that the preferred region identified in this reduced Higgs-likelihood treatment remains broadly consistent with the more detailed channel-based evaluation provided by HiggsSignals. The left panel of Fig. 6 shows the relation between $\chi^{2}_{\kappa}$ and $\chi^{2}_{\rm HS}$, with the color scale indicating the total profile-likelihood ratio.
Figure 6: Similar to Fig. 1. Left: relation between $\chi^{2}_{\kappa}$ and $\chi^{2}_{\rm HS}$, where $\chi^{2}_{\kappa}$ is derived from the $(\kappa_{V},\kappa_{F})$ plane analysis and $\chi^{2}_{\rm HS}$ is computed by the HiggsSignals package. Right: profile-likelihood distribution projected onto the signal strengths in the $\gamma\gamma$ and $b\bar{b}$ channels.
One can clearly see that the parameter points with small $\chi^{2}_{\kappa}$ also tend to yield comparatively small $\chi^{2}_{\rm HS}$ values. This result demonstrates that the simplified $(\kappa_{V},\kappa_{F})$ fit effectively captures the primary Higgs-coupling constraints relevant to our analysis, while HiggsSignals provides a more comprehensive channel-level evaluation. The spread visible in Fig. 6 indicates that the two approaches are not identical, as anticipated. In particular, the $(\kappa_{V},\kappa_{F})$ fit is based on a compressed description of the 125 GeV Higgs data and therefore does not retain the full channel-by-channel information, production-mode decomposition, or experimental correlations that are incorporated in HiggsSignals. For this reason, the $\kappa$-plane treatment should be regarded as an effective approximation rather than a complete replacement of the full Higgs likelihood. In practice, it is also somewhat more conservative than the full channel-based HiggsSignals evaluation, since the compressed $(\kappa_{V},\kappa_{F})$ description retains less experimental information and therefore has weaker discriminating power in excluding parameter points. Nevertheless, this comparison tests the robustness of our results. The conclusion that the $95~{\rm GeV}$ $\tau\tau$ signal intensity is constrained by $125~{\rm GeV}$ Higgs boson data holds true regardless of whether we use the simplified $\kappa$-plane method or the more detailed channel-based evaluation method.
The right panel of Fig. 6 shows a related feature of the surviving parameter space in the $b\bar{b}$ channel. Although the LEP $b\bar{b}$ excess is not included as an input constraint in our fit, some viable parameter points remain compatible with the corresponding preferred signal-strength range. In the GM model, the $b\bar{b}$ and $\tau\tau$ signal strengths are not independent, since both are controlled by the same underlying singlet–doublet mixing structure and share the common factor $|\kappa_{F}^{H}|^{2}\cdot\Gamma_{h}^{\rm SM}/\Gamma_{H}^{\rm tot}$, as discussed in Sec. 3. As a result, once the preferred region is selected by the combined $\gamma\gamma$ and $\tau\tau$ fit, the corresponding prediction for $\mu_{b\bar{b}}$ is automatically restricted to a relatively narrow range.
In particular, the best-fit region with $\mu_{\tau\tau}\sim 0.45$ tends to predict a $b\bar{b}$ signal strength close to the central value of the LEP excess. We do not interpret this as a primary fitting target of the present work, but rather as a correlated outcome of the surviving region selected by the LHC-driven channels. This behavior can be understood qualitatively from the fact that, unlike the $\tau\tau$ excess, the $b\bar{b}$ signal strength does not encounter the same direct tension with the precision data of the 125 GeV Higgs boson. Consequently, once the Higgs-coupling constraints compress the viable parameter space, the corresponding prediction in the $b\bar{b}$ channel can still remain close to the LEP preferred range.
Our result should not be viewed as being in conflict with earlier studies, but rather as a complementary perspective. Ref. Chen:2023bqr employed a more complete treatment of the Higgs-sector constraints and concluded that the GM model can account for the diphoton excess and remain compatible with the LEP $b\bar{b}$ excess, but cannot reproduce the large central value of the CMS $\tau\tau$ excess. Ref. Ahriche:2023wkj further showed that a larger $\tau\tau$ rate can arise either in a tiny single-peak region or, more efficiently, in a twin-peak setup involving multiple nearly degenerate states. In contrast, our analysis focuses on a single light $CP$-even state and shows that, within this setup, the $\tau\tau$ signal strength can be pushed to a moderate level of order $0.5$, while larger values are limited by the 125 GeV Higgs data. In this sense, the $b\bar{b}$ result shown in the right panel of Fig. 6 should be understood as a correlated implication of the surviving parameter space, rather than as evidence for a fully unified explanation of all three excesses.